# Relative uniform convergence and Archimedean property in pre-ordered vector spaces

### v1

@@ 1. $i)$
A quotient map $\phi_{0}:X\to(X/A,[W])$ is positive. By Theorem 3.3, $\left\langle(X/A,[W]),\phi_{0}\right\rangle$ is an object of $\mathcal{C}_{Arch}(X)$. We need to show that $\left\langle(X/A,[W]),\phi_{0}\right\rangle$ is initial. To this end, take an arbitrary object $\left\langle Y,\phi\right\rangle$ of $\mathcal{C}_{Arch}(X)$. It remains to find a positive operator $T:(X/A,[W])\to Y$ such that $T\circ\phi_{0}=\phi$. Since $Y$ is Archimedean, it follows from Lemma 3.3 that $Y_{+}$ is ru-closed, and hence is $\tau_{ru}$-closed. First, we show $\phi(W)\subseteq Y_{+}$. Let $x\in W$. There exists a net $(x_{\alpha})$ such that $X_{+}\ni x_{\alpha}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}x$. By Lemma 3.2, the operator $\phi:X\to Y$ is $\tau_{ru}$-continuous. Therefore, we obtain $Y_{+}\ni\phi(x_{\alpha})\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}\phi(x)$. As $Y_{+}$ is $\tau_{ru}$-closed, $\phi(x)\in Y_{+}$. Thus, $\phi(W)\subseteq Y_{+}$ and $\phi(-W)\subseteq-Y_{+}$, and hence $\phi(A)\subseteq Y_{+}\bigcap(-Y_{+})=\{0\}$. So, $T:(X/A,[W])\to Y$ such that $T[x]:=\phi(x)$ is well defined and positive. Clearly, $T\circ\phi_{0}=\phi$. ∎
For applications of Theorem 3.5, it is useful to describe constructively the $\tau_{ru}$-closure of the positive wedge $X_{+}$ of a POVS $X$. The idea comes from [10], where the pseudo uniform closure of a subset of a vector lattice was studied.
Adopting the idea to the POVS setting, for a subset $S$ of a POVS $X$ we denote by $S^{\prime}_{ru}$ the set of all $x\in X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ for some sequence $(x_{n})_{n}$ in $S$. Denote by $S^{(\alpha)}_{ru}$ the set $S^{(\alpha-1)}_{ru}$ when the ordinal $\alpha$ is not limit, otherwise $S^{(\alpha)}_{ru}:=\bigcup\limits_{\beta\in\alpha}S^{(\beta)}_{ru}$. It is straightforward to see that $S^{(\alpha_{1})}_{ru}\subseteq S^{(\alpha_{2})}_{ru}$ whenever $\alpha_{1}\leqslant\alpha_{2}$, and $\overline{S}_{ru}=\bigcup\limits_{\alpha\in\omega_{1}}S^{(\alpha)}_{ru}$, where $\omega_{1}$ is the first uncountable ordinal (cf., [10, Theorem 3.3] in the vector lattice setting). In particular, we have $\overline{S}_{ru}=S^{\prime}_{ru}\ \Longleftrightarrow\ S^{\prime}_{ru}=(S^{\prime}_{ru})^{\prime}_{ru}$. So, in the notations of Theorem 3.5, $W=\overline{(X_{+})}_{ru}=\bigcup\limits_{\alpha\in\omega_{1}}(X_{+})^{(\alpha)}_{ru}$. Note that each $(X_{+})^{(\alpha)}_{ru}$ is a wedge in $X$.
The statement of Theorem 3.5 motivates the following notion.
@@
It can be seen easily:
@@
Let us consider the following order ideals in a POVS $X$: the set $I_{0}(X):=\{0\}$ of all infinitesimals of $X$ of order 0; the set:
$I_{1}(X)=I(X):=\bigl\{x\in X:(\exists y\in X)(\forall n\in\mathbb{N})\pm x\leqslant\frac{1}{n}y\bigl\}$
of all infinitesimals of order 1 (or, just infinitesimals of $X$); and the set
@@
It can be seen that $I_{\alpha_{1}}(X)\subseteq I_{\alpha_{2}}(X)\subseteq(X_{+})^{(\alpha_{2})}_{ru}$ for $\alpha_{1}\leqslant\alpha_{2}$. Clearly, a POVS $X$ is almost Archimedean iff $I(X)=\{0\}$. If a POVS $X$ has a nonzero order unit $u\in X$ (i.e., $u$ satisfies $X=\bigcup\limits_{n\in\mathbb{N}}[-nu,nu]$) then $u\notin I(X)$.

### v3

@@ 4. $iv)$
A quotient map $\phi_{0}:X\to(X/A,[W])$ is positive. By Theorem 3.3, $\left\langle(X/A,[W]),\phi_{0}\right\rangle$ is an object of $\mathcal{C}_{Arch}(X)$. We need to show that $\left\langle(X/A,[W]),\phi_{0}\right\rangle$ is initial. To this end, take an arbitrary object $\left\langle Y,\phi\right\rangle$ of $\mathcal{C}_{Arch}(X)$. It remains to find a positive operator $T:(X/A,[W])\to Y$ such that $T\circ\phi_{0}=\phi$. Since $Y$ is Archimedean, it follows from Proposition 3.1 that $Y_{+}$ is ru-closed, and hence is $\tau_{ru}$-closed. First, we show $\phi(W)\subseteq Y_{+}$. Let $x\in W$. There exists a net $(x_{\alpha})$ such that $X_{+}\ni x_{\alpha}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}x$. By Lemma 3.4, the operator $\phi:X\to Y$ is $\tau_{ru}$-continuous. Therefore, we obtain $Y_{+}\ni\phi(x_{\alpha})\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}\phi(x)$. As $Y_{+}$ is $\tau_{ru}$-closed, $\phi(x)\in Y_{+}$. Thus, $\phi(W)\subseteq Y_{+}$ and $\phi(-W)\subseteq-Y_{+}$, and hence $\phi(A)\subseteq Y_{+}\bigcap-Y_{+}=\{0\}$. So, $T:(X/A,[W])\to Y$ such that $T[x]:=\phi(x)$ is well defined and positive. Clearly, $T\circ\phi_{0}=\phi$. ∎
Lemma 3.3 implies (in the notations of Theorem 3.4) that $W=\overline{(X_{+})}_{ru}=(X_{+})^{(\omega_{1})}_{ru}$. Theorem 3.4 motivates the following.
@@
Let $X$ be a POVS. Following [3], denote $D_{X}:=\{x\in X|(\exists\xi\in X_{+})(\forall n\in\mathbb{N})\ x+n^{-1}\xi\in X_{+}\}$, and $N_{X}:=\bigl\{x\in X:(\exists y\in X)(\forall n\in\mathbb{N})\pm x\leqslant\frac{1}{n}y\bigl\}$ the set of all infinitesimals of $X$. For every $x,y\in X$ denote $\Omega^{x}_{y}:=\{\alpha\in\mathbb{R}:x+\alpha y\in X_{+}\}$.
Lemma 3.5.
Let $X$ be a POVS and $y\in X$. The following assertions are equivalent:
  1. $i)$
    $\Omega^{x}_{y}\neq\{0\}$ for some $x\in X_{+}$.
  2. ii)ii)
    $y\in X_{+}-X_{+}$.
Proof.
$i)\Longrightarrow ii)$ Let $x\in X_{+}$ satisfy $\Omega^{x}_{y}\neq\{0\}$. Then $x+\alpha y\in X_{+}$ for some $\alpha\neq 0$. Thus, $\alpha y=(x+\alpha y)-x\in X_{+}-X_{+}$, and hence $y=\alpha^{-1}(x+\alpha y)-\alpha^{-1}x\in X_{+}-X_{+}$.
$ii)\Longrightarrow i)$ Let $y\in X_{+}-X_{+}$. Then $y=x^{\prime}-x^{\prime\prime}$ for some $x^{\prime},x^{\prime\prime}\in X_{+}$. It follows $y\geqslant-x^{\prime\prime}$, and $x^{\prime\prime}+\alpha y\geqslant x^{\prime\prime}-\alpha x^{\prime\prime}\geqslant 0$ for all $\alpha\in[0,1]$, and hence $0,1]\subseteq\Omega^{x^{\prime\prime}}_{y}$. ∎
Obviously, $X_{+}\subseteq D_{X}=\{x\in X|(\exists\xi\in X_{+})(\forall r>0)\ x+r\xi\in X_{+}\}$. In particular, for every $x\in D_{X}$ there exists $\xi\in X_{+}$ such that $\mathbb{R}_{+}\subseteq\Omega^{x}_{\xi}$. It is clear that $D_{X}$ is a wedge.
Proposition 3.5.
Let $X$ be a POVS. Then the following holds:
  1. $i)$
    $D_{X}=(X_{+})^{(1)}_{ru}$
  2. ii)ii)
    $D_{X}\cap-D_{X}=N_{X}$.
  3. $iii)$
    If $X$ is almost Archimedean then $D_{X}$ is a cone.
Proof.
$i)$ If $x\in D_{X}$ then, for some $\xi\in X_{+}$ we have $x+n^{-1}\xi\in X_{+}$ for all $n\in\mathbb{N}$. Since $x+n^{-1}\xi\operatorname{\xrightarrow[]{{\it ru}}}x$, then $x\in(X_{+})^{(1)}_{ru}$. Thus, $D_{X}\subseteq(X_{+})^{(1)}_{ru}$.
If $x\in(X_{+})^{(1)}_{ru}$, there exist $\xi\in X_{+}$ and a sequence $(x_{n})$ in $X_{+}$ satisfying $\pm(x_{n}-x)\leqslant n^{-1}\xi$ for all $n\in\mathbb{N}$. Then $0\leqslant x_{n}\leqslant x+n^{-1}\xi$, and hence $x+n^{-1}\xi\in X_{+}$ for all $n\in\mathbb{N}$. So, $x\in D_{X}$. Thus, $(X_{+})^{(1)}_{ru}\subseteq D_{X}$.
ii)ii) Clearly, $N_{X}$ is an order ideal in $X$ and $N_{X}\subseteq(X_{+})^{(1)}_{ru}$. Thus, $i)$ implies $N_{X}\subseteq D_{X}\cap-D_{X}$.
Suppose $x\in D_{X}\cap-D_{X}$. Then, for some $\xi_{1},\xi_{2}\in X_{+}$ we have
$x+r\xi_{1}\subseteq X_{+}\ \ \text{\rm and}\ \ -x+r\xi_{2}\subseteq X_{+}\ \ \ \ (\forall r>0).$
It follows $r(\xi_{1}+\xi_{2})\leqslant x\leqslant r(\xi_{1}+\xi_{2})$ for all $r>0$, and hence $x\in N_{X}$. So, $D_{X}\cap-D_{X}\subseteq N_{X}$.
$iii)$ Since $X$ is almost Archimedean then $N_{X}=\{0\}$, and hence $D_{X}\cap-D_{X}=\{0\}$ by ii)ii). ∎
In general, $D_{X}$ is not necessarily a cone even if $\text{\rm dim}(X)=2$ (take, for example, $\mathbb{R}^{2}_{lex}$).
It can be seen:
@@ 3. $iii)$
Let us consider the following order ideals in a POVS $X$: the set $I_{0}(X):=\{0\}$ of all infinitesimals of $X$ of order 0; the set: $I_{1}(X):=N_{X}$ of all infinitesimals of $X$); and the set
@@
It can be seen that $I_{\alpha_{1}}(X)\subseteq I_{\alpha_{2}}(X)\subseteq(X_{+})^{(\alpha_{2})}_{ru}$ for $\alpha_{1}\leqslant\alpha_{2}$. If a POVS $X$ has a nonzero order unit $u\in X$ (i.e., $u$ satisfies $X=\bigcup\limits_{n\in\mathbb{N}}[-nu,nu]$) then $u\notin N_{X}$.